OpenStudy (anonymous):

Solve x – 2y = 6 for y.

OpenStudy (anonymous):

x - 6 = 2y (x - 6)/2 = y (x/2) - 3 = y

OpenStudy (bahrom7893):

grace try the next one on ur own and mkuehn please stop just posting the answers, even if u have steps. Just tell them what to do and let them do it.

OpenStudy (anonymous):

I don't just post the answers. Check some of my answers.

OpenStudy (anonymous):

I'll try.

OpenStudy (bahrom7893):

okay mkuehn but this is a simple question, not here..